\begin{thebibliography}{26}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Beck and Teboulle(2004)]{beck2004conditional}
Amir Beck and Marc Teboulle.
\newblock A conditional gradient method with linear rate of convergence for
  solving convex linear systems.
\newblock \emph{Mathematical Methods of Operations Research}, 59\penalty0
  (2):\penalty0 235--247, 2004.

\bibitem[Bolte et~al.(2017)Bolte, Nguyen, Peypouquet, and
  Suter]{bolte2017error}
J{\'e}r{\^o}me Bolte, Trong~Phong Nguyen, Juan Peypouquet, and Bruce~W Suter.
\newblock From error bounds to the complexity of first-order descent methods
  for convex functions.
\newblock \emph{Mathematical Programming}, 165\penalty0 (2):\penalty0 471--507,
  2017.

\bibitem[Clarkson(2008)]{clarkson2008coresets}
Kenneth~L Clarkson.
\newblock Coresets, sparse greedy approximation, and the frank-wolfe algorithm.
\newblock In \emph{Proceedings of the nineteenth annual ACM-SIAM symposium on
  Discrete algorithms (SODA)}, pages 922--931. Society for Industrial and
  Applied Mathematics, 2008.

\bibitem[Demyanov and Rubinov(1970)]{demyanov1970approximate}
Vladimir~Fedorovich Demyanov and Aleksandr~Moiseevich Rubinov.
\newblock \emph{Approximate methods in optimization problems}, volume~32.
\newblock Elsevier Publishing Company, 1970.

\bibitem[Dunn(1979)]{dunn1979rates}
Joseph~C Dunn.
\newblock Rates of convergence for conditional gradient algorithms near
  singular and nonsingular extremals.
\newblock \emph{SIAM Journal on Control and Optimization}, 17\penalty0
  (2):\penalty0 187--211, 1979.

\bibitem[Frank and Wolfe(1956)]{frank1956algorithm}
Marguerite Frank and Philip Wolfe.
\newblock An algorithm for quadratic programming.
\newblock \emph{Naval research logistics quarterly}, 3\penalty0 (1-2):\penalty0
  95--110, 1956.

\bibitem[Freund and Grigas(2016)]{freund2016new}
Robert~M Freund and Paul Grigas.
\newblock New analysis and results for the frank--wolfe method.
\newblock \emph{Mathematical Programming}, 155\penalty0 (1-2):\penalty0
  199--230, 2016.

\bibitem[Garber and Hazan(2013)]{garber2013playing}
Dan Garber and Elad Hazan.
\newblock Playing non-linear games with linear oracles.
\newblock In \emph{Foundations of Computer Science (FOCS), 2013 IEEE 54th
  Annual Symposium on}, pages 420--428, 2013.

\bibitem[Garber and Hazan(2015)]{pmlr-v37-garbera15}
Dan Garber and Elad Hazan.
\newblock Faster rates for the frank-wolfe method over strongly-convex sets.
\newblock In \emph{Proceedings of the 32nd International Conference on Machine
  Learning (ICML)}, pages 541--549, 2015.

\bibitem[Gu{\'e}lat and Marcotte(1986)]{guelat1986some}
Jacques Gu{\'e}lat and Patrice Marcotte.
\newblock Some comments on wolfe's Ôaway stepÕ.
\newblock \emph{Mathematical Programming}, 35\penalty0 (1):\penalty0 110--119,
  1986.

\bibitem[Hazan(2008)]{hazan2008sparse}
Elad Hazan.
\newblock Sparse approximate solutions to semidefinite programs.
\newblock In \emph{Latin American symposium on theoretical informatics}, pages
  306--316. Springer, 2008.

\bibitem[Jaggi(2013)]{jaggi2013revisiting}
Martin Jaggi.
\newblock Revisiting frank-wolfe: projection-free sparse convex optimization.
\newblock In \emph{Proceedings of the 30th International Conference on Machine
  Learning (ICML)}, pages 427Ñ--435, 2013.

\bibitem[Lacoste-Julien and Jaggi(2013)]{lacoste2013affine}
Simon Lacoste-Julien and Martin Jaggi.
\newblock An affine invariant linear convergence analysis for frank-wolfe
  algorithms.
\newblock \emph{arXiv preprint arXiv:1312.7864}, 2013.

\bibitem[Lacoste-Julien and Jaggi(2015)]{lacoste2015global}
Simon Lacoste-Julien and Martin Jaggi.
\newblock On the global linear convergence of frank-wolfe optimization
  variants.
\newblock In \emph{Advances in Neural Information Processing Systems (NIPS)},
  pages 496--504, 2015.

\bibitem[Lan(2013)]{lan2013complexity}
Guanghui Lan.
\newblock The complexity of large-scale convex programming under a linear
  optimization oracle.
\newblock \emph{arXiv preprint arXiv:1309.5550}, 2013.

\bibitem[Lan and Zhou(2016)]{lan2016conditional}
Guanghui Lan and Yi~Zhou.
\newblock Conditional gradient sliding for convex optimization.
\newblock \emph{SIAM Journal on Optimization}, 26\penalty0 (2):\penalty0
  1379--1409, 2016.

\bibitem[Levitin and Polyak(1966)]{levitin1966constrained}
ES~Levitin and BT~Polyak.
\newblock Constrained minimization methods.
\newblock \emph{USSR Computational Mathematics and Mathematical Physics},
  6\penalty0 (5):\penalty0 1--50, 1966.

\bibitem[Li and Pong(2017)]{li2017calculus}
Guoyin Li and Ting~Kei Pong.
\newblock Calculus of the exponent of kurdyka--{\l}ojasiewicz inequality and
  its applications to linear convergence of first-order methods.
\newblock \emph{Foundations of Computational Mathematics}, pages 1--34, 2017.

\bibitem[Liu and Yang(2017)]{liu2017adaptive}
Mingrui Liu and Tianbao Yang.
\newblock Adaptive accelerated gradient converging method under h\"{o}lderian
  error bound condition.
\newblock In \emph{Advances in Neural Information Processing Systems}, pages
  3104--3114, 2017.

\bibitem[Narasimhan(2018)]{narasimhan2018learning}
Harikrishna Narasimhan.
\newblock Learning with complex loss functions and constraints.
\newblock In \emph{International Conference on Artificial Intelligence and
  Statistics}, pages 1646--1654, 2018.

\bibitem[Nesterov(2018)]{nesterov2018complexity}
Yu~Nesterov.
\newblock Complexity bounds for primal-dual methods minimizing the model of
  objective function.
\newblock \emph{Mathematical Programming}, 171\penalty0 (1-2):\penalty0
  311--330, 2018.

\bibitem[Rockafellar(1970)]{rockafellar1970convex}
R~Tyrrell Rockafellar.
\newblock \emph{Convex Analysis}.
\newblock Princeton University Press, 1970.

\bibitem[Xu et~al.(2016)Xu, Yan, Lin, and Yang]{xu2016homotopy}
Yi~Xu, Yan Yan, Qihang Lin, and Tianbao Yang.
\newblock Homotopy smoothing for non-smooth problems with lower complexity than
  $\text{O}(1/\epsilon)$.
\newblock In \emph{Advances in Neural Information Processing Systems (NIPS)},
  pages 1208--1216, 2016.

\bibitem[Xu et~al.(2017{\natexlab{a}})Xu, Lin, and Yang]{pmlr-v70-xu17a}
Yi~Xu, Qihang Lin, and Tianbao Yang.
\newblock Stochastic convex optimization: Faster local growth implies faster
  global convergence.
\newblock In \emph{Proceedings of the 34th International Conference on Machine
  Learning (ICML)}, pages 3821--3830, 2017{\natexlab{a}}.

\bibitem[Xu et~al.(2017{\natexlab{b}})Xu, Liu, Lin, and
  Yang]{DBLP:conf/nips/XuLLY17}
Yi~Xu, Mingrui Liu, Qihang Lin, and Tianbao Yang.
\newblock {ADMM} without a fixed penalty parameter: Faster convergence with new
  adaptive penalization.
\newblock In \emph{Advances in Neural Information Processing Systems 30
  (NIPS)}, pages 1267--1277, 2017{\natexlab{b}}.

\bibitem[Yang and Lin(2018)]{yang2018rsg}
Tianbao Yang and Qihang Lin.
\newblock Rsg: Beating subgradient method without smoothness and strong
  convexity.
\newblock \emph{Journal of Machine Learning Research}, 19\penalty0 (6), 2018.

\end{thebibliography}

\documentclass{colt2012} 

\usepackage{algorithm,algorithmic}
\usepackage{amssymb, multirow, paralist, color}
\newtheorem{thm}{Theorem}

\usepackage{enumitem}

\def \w {\mathbf{w}}
\def \v {\mathbf{v}}

\def \x {\mathbf{x}}

\def \x {\mathbf{x}}

\def \z {\mathbf{z}}

\def \y {\mathbf{y}}
\def \u {\mathbf{u}}

\def \y {\mathbf{y}}

\def \x {\mathbf{x}}

\def \z {\mathbf{z}}
\def \u {\mathbf{u}}

\def \w {\mathbf{w}}

\def \v {\mathbf{v}}

\begin{document}

\title[ ]{Frank-Wolfe Method is Automatically Adaptive to Error Bound Condition}
\author{\Name{Yi Xu} \Email{yi-xu@uiowa.edu}\\
\Name{Tianbao Yang} \Email{tianbao-yang@uiowa.edu}\\
\addr Department of Computer Science, The University of Iowa, Iowa City, IA 52242
}

\maketitle
\vspace*{-0.5in}
\begin{center}{October 10, 2018}\end{center}

\begin{abstract}
Error bound condition has recently gained revived interest  in optimization. It has been leveraged to derive faster convergence for many popular algorithms, including subgradient methods, proximal gradient method and accelerated proximal gradient method. However, it is still unclear whether the Frank-Wolfe (FW) method can enjoy faster convergence under error bound condition. In this short note, we give an affirmative answer to this question. We show that the FW method (with a line search for the step size) for optimization over a strongly convex set is automatically adaptive to the error bound condition of the problem. In particular, the iteration complexity  of FW can be characterized by  $O(\max(1/\epsilon^{1-\theta}, \log(1/\epsilon)))$ where $\theta\in[0,1]$ is a constant that characterizes the error bound condition. Our results imply that if the constrained set is characterized by a strongly convex function and the objective function can achieve a smaller value outside the considered domain, then the FW method enjoys a fast rate of $O(1/t^2)$. 
\end{abstract}

\section{Introduction}
In this draft, we consider the following constrained convex optimization problem:
\begin{align}\label{eqn:opt}
\min_{\x\in\Omega} f(\x)
\end{align}
where $f(\w)$ is a smooth function and $\Omega\subseteq\mathbf E$ is a bounded strongly convex set. We assume that linear optimization over $\Omega$ is much more cheaper than projection onto $\Omega$, which makes the FW method more suitable for solving the above problem than gradient methods. The goal of this paper is to show that the FW method is automatically adaptive to an error bound condition of the optimization problem. Below, we will first review the FW method and the error bound condition. In next section, we will prove that the FW method is automatically adaptive to the error bound condition. 

The original FW method, introduced by \cite{frank1956algorithm} (a.k.a. Conditional Graident method~\citep{levitin1966constrained}), is a projection-free fist-order method for minimizing smooth convex objective functions over a convex set. In recent years, the FW method has gained an increasing interest in large-scale optimization and machine learning (e.g., \citep{pmlr-v37-garbera15, freund2016new, nesterov2018complexity, narasimhan2018learning}). 	
Many existing works have shown the convergence rate of the standard FW method is $O(1/t)$ even for strongly convex objectives~\citep{clarkson2008coresets,hazan2008sparse,jaggi2013revisiting}, and in general the rate could not be improved. Under different assumptions or for some special cases, a series of works tried to get faster rates of the FW method and its variants~\citep{levitin1966constrained,demyanov1970approximate, dunn1979rates, guelat1986some,beck2004conditional,garber2013playing, lan2013complexity, lacoste2013affine, pmlr-v37-garbera15,lacoste2015global, lan2016conditional}. For example, for  minimizing smooth and strongly convex objective functions over a strongly convex set, \cite{pmlr-v37-garbera15} showed that the FM method enjoyed fast rate of $O(1/t^2)$.  

In this paper, we first consider the FW method shown in Algorithm~\ref{alg:0}, where $L_f$ denotes a smoothness constant of $f(\x)$ with respect to $\|\cdot\|$ such that $f(\x)\leq f(\y) + \nabla f(\y)^{\top}(\x - \y) + \frac{L_f}{2}\|\x - \y\|^2$ holds for any $\x, \y\in\Omega$. Note that both options for selecting the step size have been considered in the literature~\citep{jaggi2013revisiting, pmlr-v37-garbera15}. Option I requires evaluating the objective function but does not need to  know  the smoothness constant. Option II could be cheaper but requires knowing the Lipschitz constant of the gradient. Our analysis applies to both options. In the sequel, we will focus on option I, with which we have
\begin{align}\label{eqn:key}
f(\x_{t+1})&\leq f(\x_t + \eta(\y_t - \x_t)), \forall \eta\in[0,1]\nonumber\\
&\leq f(\x_t) + \eta(\y_{t} - \x_t)^{\top}\nabla f(\x_t) + \frac{\eta^2L_f}{2}\|\y_t - \x_t\|^2, \forall \eta\in[0,1]
\end{align}	
Note that for option II, the second inequality above still holds.
	
We consider the following definition of error bound condition for the optimization problem~\eqref{eqn:opt}. 
\begin{definition}[H\"{o}lderian error bound (HEB)]\label{def:1}
A function $f(\x)$ is said to satisfy a HEB condition on  $\Omega$ if there exist $\theta\in[0,1]$ and $0<c<\infty$ such that for any $\x\in\Omega$
\begin{align}\label{eqn:leb}
\min_{\w\in\Omega_*}\|\x - \w\|\leq c(f(\x) - f_*)^{\theta}.
\end{align}
where $\Omega_*$ denotes the optimal set of $\min_{\x\in\Omega}f(\x)$ and $f_*$ denotes the optimal objective value.  
\end{definition}
It is notable that $\theta=0$ is a trivial condition since it always hold due to that  $\Omega$ is a compact set.  The above HEB condition has been considered for deriving faster convergence of subgradient methods~\citep{yang2018rsg}, proximal gradient method~\citep{liu2017adaptive}, accelerated gradient method~\citep{xu2016homotopy}, and stochastic subgradient methods~\citep{pmlr-v70-xu17a}. It has been shown that many problems satisfy the above condition~\citep{xu2016homotopy, pmlr-v70-xu17a, DBLP:conf/nips/XuLLY17, liu2017adaptive, yang2018rsg}. For example, when functions are semi-algebraic and ``regular" (for instance, continuous), the above inequality is known to hold on any compact set (c.f. \citep{bolte2017error} and references therein). 
		
The last definition in this section is regarding the strongly convex set. 
\begin{definition}
A convex set $\Omega$ is a $\alpha$-strongly convex with respect to $\|\cdot\|$ if for any $\x, \y\in\Omega$, any $\gamma\in[0,1]$ and any vector $\z\in\mathbf E$ such that $\|\z\|=1$, it holds that 
\begin{align*}
\gamma\x + (1-\gamma)\y + \gamma(1-\gamma) \frac{\alpha}{2}\|\x - \y\|^2\z\in\Omega. 
\end{align*}
\end{definition}
{\bf Remark. }Many previous works~(e.g., \citep{levitin1966constrained,demyanov1970approximate, dunn1979rates, pmlr-v37-garbera15}) considered this condition of feasible set when studying the FW method.
\begin{algorithm}[t]
\caption{Frank-Wolfe Method} \label{alg:0}
\begin{algorithmic}
\STATE Initilization: $\x_0 \in\Omega$
\FOR{$t=0,\ldots, T$}
    	\STATE Compute $\y_t \in \arg\min_{\y\in\Omega}\nabla f(\x_t)^{\top}\y$\;
	\STATE Option I: Set $\eta_t = \arg\min_{\eta\in[0,1]}f(\x_t + \eta(\y_t - \x_t))$ \;
	\STATE Option II: Set $\eta_t = \arg\min_{\eta\in[0,1]}\eta(\y_t - \x_t)^{\top}\nabla f(\x_t) + \frac{\eta^2L_f}{2}\|\y_t - \x_t\|^2$ \;
 	\STATE Compute $\x_{t+1} =\x_t + \eta_t (\y_t - \x_t) $\;
\ENDFOR
\end{algorithmic}
\end{algorithm}

\section{Adaptive Convergence of the FW method}
In this section, we show that the FW method is {\it automatically} adaptive to the HEB condition, enjoying a faster convergence rate than the standard $O(1/t)$ rate without the knowledge of the HEB condition. 

We first prove the following lemma. 
\begin{lemma}\label{lem:1}
Assume $f(\x)$ obeys the HEB condition on $\Omega$ with $\theta\in[0,1]$, then it holds that 
\begin{align*}
\|\nabla f(\x)\|_*\geq \frac{1}{c}(f(\x) - f_*)^{1-\theta}.
\end{align*}
\end{lemma}
\begin{proof}
Let $\x_*$ denote the optimal solution in $\Omega_*$ that is closest to $\x$ measured in $\|\cdot\|$.  
By convexity of $f(\cdot)$, we have
\begin{align*}
f(\x_*)\geq f(\x) + \nabla f(\x)^{\top}(\x_* - \x).
\end{align*}
Thus, 
\begin{align*}
f(\x) - f(\x_*)\leq \|\nabla f(\x)\|_* \|\x - \x_*\|\leq c(f(\x) - f_*)^{\theta} \|\nabla f(\x)\|_* .
\end{align*}
As a result, 
\begin{align*}
\|\nabla f(\x)\|_*\geq \frac{1}{c}(f(\x) - f_*)^{1-\theta}.
\end{align*}
\end{proof}
The second lemma is from \citep{pmlr-v37-garbera15}.
\begin{lemma}\label{lem:2}
For the FW method given in Algorithm~\ref{alg:0}, for $t=0,\ldots, $ we have
\begin{align*}
f(\x_{t+1}) - f_*\leq (f(\x_t) - f_*)\max\left\{\frac{1}{2}, \left(1 - \frac{\alpha \|\nabla f(\x_t)\|_*}{8L_f}\right)\right\}.
\end{align*}
\end{lemma}
Finally, we prove the following theorem. 
\begin{thm}
For every $t\geq 1$, we have
\begin{align*}
f(\x_t) -f _* \leq \left\{\begin{array}{lc}\frac{C}{(t+k)^{1/(1-\theta)}}& \text{ if }\theta\in [0,1)\\ \\  \rho^{t}(f(\x_0) - f_*)& \text{ otherwise}\end{array}\right.
\end{align*}
where $k\geq \max\left\{\frac{2 - 2^{1-\theta}}{2^{1-\theta} -1}, C'\right\}$,  $C\geq \max\left\{\frac{L_fD^2(1+k)^{\frac{1}{1-\theta}}}{2}, 2\left(\frac{C'}{M}\right)^{\frac{1}{1-\theta}}\right\}$, $C' = \frac{1}{1-\theta -\theta(2^{1-\theta}-1)}$, and $\rho=\max\left\{\frac{1}{2}, 1 - \frac{\alpha}{8cL_f}\right\}$. 
\end{thm}
{\bf Remark. }In order to find an $\epsilon$-approximate solution $\x_t$ such that $f(\x_t) - f_* \leq \epsilon$, the iteration complexity of FW method is $O(\max(1/\epsilon^{1-\theta}, \log(1/\epsilon)))$ with $\theta\in[0,1]$.
\begin{proof}
When $\theta=1$, the conclusion is trivial, which follows directly from Lemma~\ref{lem:2}. Next, we prove for $\theta\in[0,1)$. 
Let $\beta=1-\theta$. $h_t = f(\x_t) - f_*$. Combining Lemma~\ref{lem:1} and Lemma~\ref{lem:2}, we have
\begin{align}\label{eqn:max}
h_{t+1} \leq h_t \max\left\{\frac{1}{2}, 1 - \frac{\alpha}{8cL_f}h_t^{\beta}\right\} =  h_t \max\left\{\frac{1}{2}, 1 - M h_t^{\beta}\right\}
\end{align}

We prove by induction that $h_t\leq \frac{C}{(t+k)^{1/\beta}}$. 

For the case $t=1$, following~(\ref{eqn:key}) we have
\begin{align*}
h_1\leq  h_0(1-\eta) + \frac{L_f\eta^2D^2}{2} \leq \max\left\{ \frac{L_fD^2}{2}, h_0 \right\} \leq \frac{L_fD^2}{2}, \forall \eta\in[0,1],
\end{align*}
where we use the fact that $h_0 = f(\x_0) - f(\x_*) \le \nabla f(\x_*)^\top(\x_0 - \x_*) + \frac{L_f}{2}\|\x_0-\x_*\|^2 =  \frac{L_f}{2}\|\x_0-\x_*\|^2 \leq  \frac{L_fD^2}{2}$].
As long as $C/(1+k)^{1/\beta}\geq L_fD^2/2$, we have the conclusion holds for $t=1$.

Next, we consider $t\geq 1$. First assume that the max operation in~(\ref{eqn:max}) gives $1/2$, i.e., 
\begin{align*}
h_{t+1}\leq \frac{h_t }{2}\leq \frac{C}{2(t+k)^{1/\beta}}\leq \frac{C}{(t+1+k)^{1/\beta}}\frac{(t+1+k)^{1/\beta}}{2(t+k)^{1/\beta}}\leq \frac{C}{(t+1+k)^{1/\beta}},
\end{align*}
where the last inequality holds as long as 
\begin{align*}
\frac{(t+1+k)^{1/\beta}}{2(t+k)^{1/\beta}}\leq 1, \forall t\geq 1, \quad\text{i.e.,}\quad 1 + \frac{1}{t+k}\leq 2^\beta, \forall t\geq 1, \quad\text{i.e.,}\quad  k\geq \frac{2 - 2^{\beta}}{2^\beta -1}.
\end{align*}
Next, consider the case that the max operation is the second argument. In this case, if $h_t\leq\frac{C}{2(t+k)^{1/\beta}}$, the same conclusion holds under the above condition of $k$. Otherwise, $h_t\geq \frac{C}{2(t+k)^{1/\beta}}$. We have
\begin{align*}
h_{t+1}\leq h_t (1-Mh_t^{\beta})&\leq \frac{C}{(t+k)^{1/\beta}}\left(1 - M\left(\frac{C}{2}\right)^{\beta}\frac{1}{t + k} \right)\\
&\leq \frac{C}{(t+k+1)^{1/\beta}}\frac{(t+k+1)^{1/\beta}}{(t+k)^{1/\beta}}\left(1 - M\left(\frac{C}{2}\right)^{\beta}\frac{1}{t + k} \right)\\
&\leq  \frac{C}{(t+k+1)^{1/\beta}}\left(1 +\frac{C'}{t+k}\right)\left(1 - \frac{C'}{t+k}\right)
\end{align*}
To show the last inequality holds, we can set $C' = \frac{1}{\beta -(1-\beta)(2^{\beta}-1)} > 1$ and $C\geq 2(C'/M)^{1/\beta}$. To see this, we need to show that 
\begin{align*}
\log(1+ C'x) - \frac{1}{\beta}\log(1+x)\geq 0, \forall 0\leq x\leq 2^{\beta}-1. 
\end{align*}
In fact, due to $\frac{C'}{1+C'x} - \frac{1}{\beta(1+x)}\geq 0, \forall 0\leq x\leq 2^\beta -1$, 
it gives $1 + C'x\geq (1+x)^{1/\beta}$ holds for all $0\leq x\leq 2^\beta -1$. Plugging $x=1/(t+k)\leq 2^\beta -1$ into this inequality, we get what we want $1+C'/(t+k)\geq (1+\frac{1}{t+k})^{1/\beta}$.
\end{proof}

\section{Examples}
Lastly, we give examples exhibiting the HEB condition with $\theta=1/2$. In particular, let us 
 consider 
 \begin{align}\label{ex:1}
 \min_{g(\x)\leq r} f(\x)
 \end{align}
 where $g(\x)$ is a non-negative, strongly and smooth function. It is shown that $\Omega=\{\x: g(\x)\leq r\}$ is a strongly convex set~\citep{pmlr-v37-garbera15}. 
 \begin{lemma}
 Assume that $\min_{\x}f(\x)<\min_{g(\x)\leq r}f(\x)$ and there exists a $\x_0$ such that $g(\x_0) < r$, then the above problem satisfies HEB with $\theta=1/2$.
 \end{lemma}
 \begin{proof}
 We set $\Omega = \{\x: g(\x) \leq r \}$ and $\Omega_* = \arg\min_{g(\x)\leq r} f(\x)$, and we define an indicator function as follows,
 \begin{align*}
 I_{\Omega}(\x) = 
 \left\{ \begin{array}{rl}
         0 & \text{if~}\x\in\Omega, \\ 
        +\infty & \text{if~}\x\notin\Omega.
 \end{array}\right.
 \end{align*}
Then the problem of (\ref{ex:1}) can be written as
 \begin{align*}
 \min_{\x} \widehat f(\x) := f(\x) + I_{\Omega}(\x),
 \end{align*}
 and thus we also have $\Omega_* = \arg\min_{\x}\widehat f(\x)$.
 We only need to consider any fixed $\x_*\in\Omega_*$. By the condition of $g(\x_0) < r$ and Corollary 28.2.1 of~\citep{rockafellar1970convex}, there exists $\lambda^*\geq 0$ such that
 \begin{align}\label{ex1:ineq:1}
 \nonumber \widehat f(\x_*) = & \min_{\x} \widehat f(\x) = f(\x_*) = \min_{\x\in\Omega} f(\x) = \min_{\x} \{f(\x)+\lambda^*(g(\x)-r)\}\\
 \leq & f(\x_*)+\lambda^*(g(\x_*)-r) \leq f(\x_*),
 \end{align}
 where the first inequality is due to $\x_*\in\Omega_*$; the second inequality uses the fact that $\x_*\in\Omega_*$ hence $g(\x_*)-r\leq 0$.
 Then, equality holds for (\ref{ex1:ineq:1}), which implies $f(\x_*)+\lambda^*(g(\x_*)-r) = f(\x_*)$, that is,
  \begin{align}\label{ex1:ineq:2}
\lambda^*(g(\x_*)-r) = 0.
 \end{align}
 On the other hand, let $\u_* \in \arg\min_{\x} f(\x)$,  then based on the assumption of $\min_{\x}f(\x)<\min_{g(\x)\leq r}f(\x)$ we know $\u_* \notin \Omega_*$ hence $\u_* \notin \Omega$. By (\ref{ex1:ineq:1}), we also know
 \begin{align*}
 f(\u_*) < \min_{\x\in\Omega} f(\x) =  \min_{\x} \{f(\x)+\lambda^*(g(\x)-r)\} \leq f(\u_*)+\lambda^*(g(\u_*)-r),
 \end{align*}
 which implies 
  \begin{align}\label{ex1:ineq:3}
 \lambda^*(g(\u_*)-r)>0.
 \end{align}
 Since $\u_* \notin \Omega$, then $g(\u_*)-r > 0$. In order to have (\ref{ex1:ineq:3}), we need $\lambda^* > 0$. Thus, by (\ref{ex1:ineq:2}) we have
   \begin{align}\label{ex1:ineq:4}
g(\x_*)-r = 0.
 \end{align}
 For any such $\lambda^*>0$, then by Theorem 28.1 of~\citep{rockafellar1970convex}, we also have
 \begin{align}\label{ex1:ineq:5}
 \Omega_* = \{\x: g(\x) = r\} \cap \arg\min_{\x}\{f(\x) + \lambda^*(g(\x)-r)\}.
 \end{align}
 Since $g(\x)$ is strongly convex, $f(\x)$ is convex and $\lambda^*>0$, then $f(\x) + \lambda^*(g(\x)-r)$ is also strongly convex, implying that 
 $\v_* = \arg\min_{\x}\{f(\x) + \lambda^*(g(\x)-r)\}$ is a unique constant.
 Due to $\lambda^*>0$, $g(\v_*)$ is also a constant~\citep{li2017calculus}.
 By (\ref{ex1:ineq:5}) we have $g(\v_*) = g(\x_*) = r$.
 Therefore,
  \begin{align}\label{ex1:ineq:6}
 \Omega_* =  \arg\min_{\x}\{f(\x) + \lambda^*(g(\x)-r)\}.
 \end{align}
 By the strong convexity of $f(\x) + \lambda^*(g(\x)-r)$ we know for any $\x\in\Omega$ and $\x_*\in\Omega_*\subseteq \Omega$,
\begin{align*}
\frac{1}{c^2}\|\x-\x_*\|^2 \leq f(\x) + \lambda^*(g(\x)-r) - [f(\x_*) + \lambda^*(g(\x_*)-r)],
\end{align*}
where $c>0$.
Since $\lambda^*>0$, $g(\x)-r \leq 0$ and $g(\x_*)-r=0$, we get
 \begin{align*}
\frac{1}{c^2}\|\x-\x_*\|^2 \leq f(\x)  - f(\x_*).
\end{align*}
Therefore, for any $\x\in\Omega$
\begin{align*}
\min_{\w\in\Omega_*} \|\x - \w\| \leq c(f(\x) - f_*)^{1/2},
\end{align*}
which implies $\theta = 1/2$.
 \end{proof}

\bibliography{all}

\end{document}